OpenStudy (anonymous):

how would you find the solutions for: e^ { 2 x } - 6 e^ { x } +8 = 0

myininaya (myininaya):

let u=e^x then factor

OpenStudy (anonymous):

what do you mean by u

myininaya (myininaya):

let u=e^x

myininaya (myininaya):

e^x is u replace e^x with u replace e^(2x) with u^2

myininaya (myininaya):

then factor the quadratic

myininaya (myininaya):

then replace u with e^x

OpenStudy (anonymous):

i dont think im deriving the solutions

myininaya (myininaya):

\[u^2-6u+8=0\] so you didn't replace e^x with u? you just factor now (u-2)(u-4)=0 u=2 or u=4 but u=e^x so replace u with e^x e^x=2 or e^x=4 then solve for x